Suppose that at a busy traffic junction, the probability p of an individual car having
an accident is 0.0001. During a certain part of the day, 100 cars pass through the
junction. What is the probability of two or more cars being involved in an accident
within this period?
1
Expert's answer
2020-06-07T13:37:19-0400

"\lambda=0.0001\times 100=0.01\\
P(X\geq 2)=1-P(X<2)\\
=1-[P(X=0)+P(X=1)]\\
=1-[\frac{e^{-0.01}(0.01)^{0}}{0!}+\frac{e^{-0.01}(0.01)^{1}}{1!}]\\
=1-0.99995=0.00005"